OpenStudy (anonymous):

One saline solution contains .75 g NaCl in exactly 100 mL of solution. What is the molarity of the solution?

OpenStudy (anonymous):

How many moles in a gram?

OpenStudy (anonymous):

it depends on compound and molecular weight. Na=23g and Cl=35g?

OpenStudy (anonymous):

Right, so NaCl has a molar mass of about 58 g/mol, or approximately .0172 mol/g \[\frac{.75\text{ g}}{100\text{ mL}}\times\frac{.0172\text{ mol}}{1\text{ g}}\times\frac{1000\text{ mL}}{1\text{ L}}=\cdots\frac{\text{mol}}{\text{L}}\]

OpenStudy (anonymous):

could you show me how to cancel each one out I don't know how to

OpenStudy (anonymous):

The units cancel: \[\frac{.75\color{blue}{\cancel{\text{ g}}}}{100\color{red}{\cancel{\text{ mL}}}}\times\frac{.0172\text{ mol}}{1\color{blue}{\cancel{\text{ g}}}}\times\frac{1000\color{red}{\cancel{\text{ mL}}}}{1\text{ L}}=\cdots\frac{\text{mol}}{\text{L}}\] The rest of the work is just computation. .75 divided by 100 times .0172 times 1000 is ...

OpenStudy (anonymous):

yep, yw

OpenStudy (anonymous):

It should be, yes. Don't forget your units. It might not be the exact answer due to rounding.